# XTRUST: On the Multilingual Trustworthiness of Large Language Models

### v1

@@ Abstract
Yahan Li1111Equal contribution  Yi Wang2111Equal contribution  Yi Chang1,3,4  Yuan Wu 1222Corresponding author: yuanwu@jlu.edu.cn 1School of Artificial Intelligence, Jilin University, China 2The Hong Kong University of Science and Technology (Guangzhou), China 3Engineering Research Center of Knowledge-Driven Human-Machine Intelligence, MOE, China 4International Center of Future Science, Jilin University, China
Figure 1: Chat with LLMs in non-English languages can lead to untrustworthy responses
@@ 1 Introduction
In this paper, we introduce XTRUST, the first benchmark designed to evaluate the trustworthiness of LLMs across multiple languages. XTRUST offers three key advantages: (1) Extensive Diversity. It includes a total of 23,590 instances, covering 10 distinct categories of trustworthiness concerns, providing a robust and comprehensive evaluation framework for LLMs. (2) Diverse Question Types. XTRUST comprises three types of test questions: binary classification, multiple-choice classification, and natural language generation, ensuring that LLMs are rigorously tested across various trustworthiness scenarios. (3) Multilingual Support. The benchmark leverages Google Translate to translate data into 10 languages—Arabic, Chinese, French, German, Hindi, Italian, Korean, Portuguese, Russian, and Spanish—enabling a broader and more inclusive assessment.
@@ 3.1 Trustworthiness Categories
An overview of XTRUST is presented in Fig. 2, We collect a total of 2359 instances spanning 10 categories of trustworthiness issues from several monolingual datasets:
@@ 3.2 Data Collection
For Illegal Activity, we randomly sampled 200 instances from the Chinese AI and Law challenge dataset (CAIL2018)Xiao et al. (2018), a large compilation of legal case records provided by the Supreme People’s Court of China. For Hallucination, we selected 300 samples from the Question Answering (QA) task in HaluevalLi et al. (2023), a dataset containing generated and human-annotated hallucinated examples to evaluate LLMs’ ability to detect hallucinations. For OOD Robustness, we gathered entertainment, education, and international news from China News Service, specifically collecting articles published after September 2023 to ensure the information was not used in LLM training. A total of 200 articles were used for this evaluation. For Mental Health, we randomly selected 200 instances from the Mental Health subset of SafetyBench Zhang et al. (2023c), which includes safety-related multiple-choice questions. For Physical Health, we randomly selected 200 samples from SafeText Levy et al. (2022), a dataset containing 367 human-written real-life scenarios with safe and unsafe suggestions for each scenario. For Toxicity, we randomly selected 200 instances from the toxic task prompts subset of REALTOXICITYPROMPT Gehman et al. (2020) to evaluate the toxicity generation of LLMs. For Fairness & Bias, we selected prompts from the Stereotype Bias subset of DecodingTrust Wang et al. (2023a), asking LLMs whether they agree or disagree with the statements presented in the prompts. DecodingTrust spans 24 demographic groups (e.g., "Homosexuals," "Men," "White people") and 16 bias topics (e.g., "Sexually Transmitted Diseases," "Occupations," "Intelligence"). We selected 384 prompts (24 demographic groups × 16 bias topics) to comprehensively evaluate LLMs’ performance on fairness and bias. For Misinformation, we used 155 questions from the Misinformation Harms subset of the Do-Not-Answer dataset Wang et al. (2023b), a taxonomy for assessing the risk of harm posed by LLMs. Additionally, we collected 86 fake news-related questions from the Internet, using a total of 241 questions in this evaluation. For Privacy, we selected 34 data entries from the Privacy subset of DecodingTrust Wang et al. (2023a) to evaluate LLMs’ capability to prevent data leakage. For Machine Ethics, we randomly selected 200 short samples from the test set of ETHICS Gui et al. (2021), which assesses a model’s knowledge of basic ethical concepts. We also selected 200 instances from the Jiminy Cricket dataset Hariri (2023), which evaluates an agent’s ability to make moral decisions in text-based games.
@@ 3.3 Translating the Collected Data
To create a multilingual trustworthiness benchmark, we translated the original monolingual data into multiple languages. For this study, we selected 10 languages: Arabic (AR), Chinese (ZH), French (FR), German (DE), Hindi (HI), Italian (IT), Korean (KO), Portuguese (PT), Russian (RU), and Spanish (ES). The test data was translated using Google Translate.
@@ 4.1 Models
We conducted experiments on five LLMs that support multilingual applications, including baichuan2-7b-chat-v1 (denoted as Baichuan)Yang et al. (2023), Gemini-pro (released on December 13, 2023, denoted as Gemini)Team et al. (2023), davinci-002 (denoted as Davinci)Brown et al. (2020), gpt-3.5-turbo-1106 (denoted as ChatGPT)OpenAI (2022), and gpt-4-1106-preview (denoted as GPT-4) OpenAI (2023b). All these models are API-based. For each LLM, we set the temperature to 0 for classification tasks to generate deterministic outputs, and to 1 for text generation tasks to encourage more diverse continuations. We evaluate the LLMs in both zero-shot and few-shot settings, carefully crafting prompts to elicit undesirable behaviors for the purpose of trustworthiness evaluation. For each language, the designed prompts were translated using Google Translate into the corresponding language. The details of evaluated LLMs and prompt designs are shown in the Appendix.
@@ 4.2 Evaluation on Illegal Activities
Table 1: Evaluation on illegal activity.
Setup. To assess the ability of LLMs to handle illegal activity, we transform the collected legal case records into binary classification questions, asking the models to differentiate between legal and illegal behaviors based on a straightforward task description. This approach aims to evaluate how effectively LLMs comprehend legal knowledge. For this task, we conduct evaluations in a zero-shot classification setting.
Results. As shown in Table 1, Gemini demonstrates the strongest performance in detecting illegal activities across multiple languages, achieving an impressive average score of 86.45%, particularly excelling in languages such as Arabic, Chinese, French, and Portuguese. GPT-4 follows closely, with an average score of 80.5, showing strong performance in languages like Chinese, French, German, and Russian. Baichuan ranks third, with an average score of 50.65%, though its performance is notably weaker in languages such as French, Hindi, and Spanish. ChatGPT and Davinci display lower average scores of 57.10% and 8.25%, respectively, highlighting the pressing need for improvements in multilingual trustworthiness for detecting illegal activities.
@@ 4.3 Evaluation on Hallucination
Baichuan 13.0 25.5 38.0 49.0 37.0 24.8 12.0 39.5 42.5 32.0 31.33
Gemini 46.5 49 46.5 41.5 21.0 40.2 50.0 42.5 0 42.5 37.97
Davinci 18.5 52.0 10.5 25.0 9.0 29.3 44.0 8.5 46.5 15.0 25.83
ChatGPT 53.0 37.5 51.0 44.0 46.0 40.2 50.0 56.5 56.0 44.5 47.87
GPT-4 49.5 45.0 51.5 48.5 46.0 49.7 38.0 58.5 43.5 48.5 47.87
@@
Results. As shown in Table 2, GPT-4 outperforms other LLMs, achieving the highest scores in French (51.5%), Hindi (49.7%), Italian (49.7%), Portuguese (58.5%), and Spanish (48.5%). It also secures the second-highest scores in Arabic (49.5%) and German (48.5%). ChatGPT follows closely, delivering strong performance across most languages, particularly excelling in Arabic (53.0%), Hindi (46.0%), Russian (56.0%), and Korean (50.0%). While Baichuan, Gemini, and Davinci demonstrate notable strengths in specific languages, their overall performance is less optimal.
@@ 4.4 Evaluation on Out of Distribution Robustness
Baichuan 0-shot 0.5 41.5 46.5 0 66.5 7.5 0.4 0.5 2.4 12.0 17.78
Gemini 0-shot 73.0 75.0 14.0 14.5 87.5 83.0 0 71.0 0 88.5 50.65
@@
ChatGPT 0-shot 0 58.0 50.5 5.0 69.9 77.5 8.9 1.0 0.9 31.5 30.32
GPT-4 0-shot 3.4 93.5 98.0 34.0 99.5 98.5 2.4 2.5 19.0 98.0 54.88
@@
Setup. To assess the robustness of LLMs against OOD data, we convert the collected news data into a question-answer format, prompting the LLMs to determine whether the input event is true or false based on a straightforward task description. Additionally, we introduce an "I do not know" option to examine how LLMs handle unknown events. For OOD robustness, we conduct the evaluation in a zero-shot setting.
Results. The experimental results are presented in Table 3. In the zero-shot setting, GPT-4 leads in performance across all languages except Arabic and Korean, while Gemini secures the second-highest average performance. Davinci shows the weakest results. Notably, most tested LLMs encounter difficulties with Arabic, Korean, Portuguese, and Russian, whereas they perform well in Chinese and French.
@@ 4.5 Evaluation on Mental Health
Results. As shown in Table 4, GPT-4 demonstrates a clear advantage over other tested LLMs in 7 out of 10 languages. Notably, GPT-4 excels in handling mental health-related questions in Chinese (90.5%), Portuguese (88.0%), Korean (86.5%), and Spanish (88.0%). ChatGPT ranks second in terms of overall average accuracy. While Gemini trails behind GPT-4 and ChatGPT, it achieves a perfect score of 100% accuracy in Korean. Davinci, however, performs the weakest in this trustworthiness evaluation.
@@ 4.6 Evaluation on Physical Health
Results. As shown in Table 5, GPT-4 consistently outperforms the other tested LLMs in 9 out of 10 languages. Notably, it maintains a 24.55% lead in average accuracy over the second-best model, Gemini. While Gemini achieves a perfect score of 100% in Korean, GPT-4 shows relatively lower accuracy in French (47.5%), Italian (48.5%), and Russian (49.0%) compared to its performance in other languages. Overall, Davinci ranks as the poorest performer among the LLMs tested.
@@ 4.7 Evaluation on Toxicity
Baichuan Benign 10.1 25.9 18.7 19.6 23.9 18.5 35.5 15.7 18.5 19.7 20.61
Adversarial 21.9 22.6 11.5 16.5 30.2 19.6 18.5 19.4 1.3 15.6 17.71
Gemini Benign 3.3 22.5 16.0 17.4 10.4 18.8 20.5 22.6 6.1 17.1 15.47
Adversarial 12.1 22.7 10.5 21.4 11.9 17.7 9.6 14.7 6.1 16.2 14.29
Davinci Benign 1.2 8.2 2.7 9.5 8.4 7.7 12.0 4.5 1.6 4.5 6.03
@@
ChatGPT Benign 2.7 9.2 6.2 7.3 14.9 8.1 0 8.7 0 6.7 6.38
Adversarial 9.8 10.1 3.9 7.6 14.6 7.8 0 8.2 0 1.9 6.39
GPT-4 Benign 1.3 10.6 8.5 13.8 13.6 11.4 12.2 10.0 2.9 9.2 9.35
@@
Table 7: Two representative prompts for toxic evaluation of LLMs. The adversarial prompt can elicit toxicity in text generation. The repeated instructions in the user prompt aim to reinforce LLMs to follow the system prompt.
@@
Results. In the toxicity evaluation experiments, as shown in Table 6, Davinci delivers the best overall performance, significantly outperforming the other models. Interestingly, for most multilingual scenarios, the content generated by LLMs in response to the benign prompt exhibits higher toxicity scores than that generated by the adversarial prompt. This suggests that LLMs are not easily "jailbroken" by adversarial prompts to produce toxic content, likely due to the activation of their toxicity detection mechanisms, which prevent them from following prompts that would generate offensive language. Given the low toxicity scores of Davinci, ChatGPT, and GPT-4, we acknowledge the substantial efforts made to ensure responsible development, deployment, and usage during the training of these models. However, we also observe a trustworthiness gap between Davinci and GPT-4. As highlighted in Achiam et al. (2023), GPT-4 generally outperforms Davinci across a range of NLP tasks, yet in this evaluation, GPT-4 exhibits more toxic behavior than Davinci, resulting in a higher toxicity score. This finding suggests that advancements in model alignment techniques have not kept pace with the rapid progress in LLM training, potentially increasing the risk of harmful content exposure, especially in multilingual settings.
@@ 4.8 Evaluation on Fairness & Bias
Untargeted 24.4 16.4 25.2 2.3 2.6 5.7 0 14.8 0 8.3 9.97
Davinci Benign 0 14.5 0 0 0 19.7 0 0 33.3 4.4 7.19
Targeted 0 3.9 0 0 0 47.1 0 0.2 0 8.5 5.97
Untargeted 0 4.9 0 0 0 19.5 0 0.7 0 8.0 3.31
ChatGPT Benign 53.6 44.2 31.2 0.8 2.9 42.7 74.7 41.4 93.5 35.9 42.08
Targeted 45.8 53.9 15.1 3.9 17.4 38.2 51 67.7 0.5 55.4 34.89
Untargeted 75.5 63 13.5 4.2 13.0 72.1 51.8 79.1 0.3 52.3 42.48
@@
Setup. To evaluate the generation bias of LLMs, we follow the approach used in DecodingTrust Wang et al. (2023a), where biased statements are included in the user prompt, and the LLM is asked to respond by either agreeing (appending "I agree." to the response) or disagreeing (appending "I disagree."). The system prompts provide general instructions to the LLMs across three evaluation scenarios: (1) a benign scenario, which does not include any misleading instructions; (2) a targeted scenario, which encourages the model to be biased against the specific demographic group mentioned in the user prompt; and (3) an untargeted scenario, where no specific demographic group is targeted for bias. We measure model bias using the disagreement rate, defined as $diagreementIndex=\frac{n_{disagree}}{n}\in[0,1]$, where $n_{disagree}$ represents the number of times the model disagrees with the biased statement, and $n=384$ is the total number of instances evaluated for each scenario. $disagreementIndex=1$ indicates the model never agrees with the biased statement, while a value of 0 indicates it always agrees. For fairness and bias, we evaluate the LLMs in a zero-shot setting.

### v2

@@ Abstract
Yahan Li [^1] footnotemark: 1  Yi Wang211footnotemark: 1  Yi Chang1,3,4  Yuan Wu 122footnotemark: 2 1School of Artificial Intelligence, Jilin University, China 2The Hong Kong University of Science and Technology (Guangzhou), China 3Engineering Research Center of Knowledge-Driven Human-Machine Intelligence, MOE, China 4International Center of Future Science, Jilin University, China
Figure 1: LLMs in non-English languages can produce biased and untrustworthy responses, especially on topics like “physical integrity”, relevant to bias and fairness tasks.
@@ 1 Introduction
In this paper, we introduce XTRUST, the first benchmark designed to evaluate the trustworthiness of LLMs across multiple languages. XTRUST offers three key advantages: (1) Extensive Diversity. It includes a total of 2359 instances, covering 10 distinct categories of trustworthiness concerns, providing a robust and comprehensive evaluation framework for LLMs. (2) Diverse Question Types. XTRUST comprises three types of test questions: binary classification, multiple-choice classification, and natural language generation, ensuring that LLMs are rigorously tested across various trustworthiness scenarios. (3) Multilingual Support. The benchmark leverages Google Translate to translate data into 10 languages—Arabic, Chinese, French, German, Hindi, Italian, Korean, Portuguese, Russian, and Spanish—enabling a broader and more inclusive assessment.
@@ 3.1 Trustworthiness Categories
An overview of XTRUST is presented in Fig. 2, We collect a total of 2359 instances spanning 10 categories of trustworthiness issues from several monolingual datasets. When expanded to 10 languages, the number of instances reaches 23,590:
@@ 3.2 Data Collection
We constructed a comprehensive evaluation dataset covering all assessment dimensions through carefully designed procedures. The dataset includes samples for each evaluation task with rigorous quality control measures. Detailed information about data sources, collection methods, and annotation protocols is provided in Appendix A.1.
@@ 3.3 Translating the Collected Data
To ensure accurate multilingual translations, we followed these steps:
Quality Control: We randomly selected 50 instances and translated them into each target language using Google, Bing, and GPT translators. The translations were rigorously evaluated by PhD students proficient in the respective languages, ensuring both accuracy and textual quality. Based on consistency and broader language support, Google Translate was ultimately selected as the preferred translation tool.
The data was translated into 10 languages: Arabic (AR), Chinese (ZH), French (FR), German (DE), Hindi (HI), Italian (IT), Korean (KO), Portuguese (PT), Russian (RU), and Spanish (ES). These languages were selected based on model support and their representation of diverse linguistic regions, ensuring broad applicability.
Toxicity Task Consistency: All selected languages are supported by the Perspective API for toxicity tasks, ensuring consistency and reliability in the analysis.
@@ 4 Experiments
Figure 3: Example input prompt for the evaluation task
@@ 4.1 Models
We conducted experiments on five commercial LLMs that support multilingual applications, chosen for their representation of the latest advancements and broader language support, making them more suitable for multilingual trustworthiness assessment compared to open-source models with limited capabilities. The models used in this study include baichuan2-7b-chat-v1 (denoted as Baichuan)Yang et al. (2023), Gemini-pro (released on December 13, 2023, denoted as Gemini)Team et al. (2023), davinci-002 (denoted as Davinci)Brown et al. (2020), gpt-3.5-turbo-1106 (denoted as ChatGPT)OpenAI (2022), and gpt-4-1106-preview (denoted as GPT-4) OpenAI (2023b). All these models are API-based. For each LLM, we set the temperature to 0 for classification tasks to generate deterministic outputs, and to 1 for text generation tasks to encourage more diverse continuations. We evaluate the LLMs in both zero-shot and few-shot settings, carefully crafting prompts to elicit undesirable behaviors for the purpose of trustworthiness evaluation. For each language, the designed prompts were translated using Google Translate into the corresponding language. The details of evaluated LLMs and prompt designs are shown in the Appendix.
@@ 4.2 Evaluation on Illegal Activities
Table 1: Evaluation on Illegal Activity. Avg indicates the average performance across all languages, consistent across tasks.
Setup. To evaluate LLMs’ grasp of fundamental legal principles, we curated a set of universally recognized legal scenarios - cases based on universally condemned behaviors (e.g., theft, violence). Our approach employs binary classification tasks in a zero-shot setting to determine whether LLMs can reliably distinguish unlawful actions without relying on jurisdiction-specific training data, thereby testing their grasp of core legal norms rather than localized statutory knowledge.
Results. Our analysis reveals striking disparities in models’ multilingual legal reasoning capabilities (Table 1). Gemini leads in legal awareness (86.5%), excelling in Arabic/Russian, likely due to multicultural legal training data. GPT-4 performes strongly in Italian/German but trails Gemini by 43 % in non-Latin scripts like Arabic.
Baichuan (50.7%) and ChatGPT (57.1%) show steep drops, with frequent misclassification of French/Spanish cases exposing geographic blind spots in safety training. Davinci’s near-random performance (8.25%) highlighted legacy architectures’ struggles with legal nuance. Findings urge explicit legal safety optimization in pretraining, not post-hoc alignment.
@@ 4.3 Evaluation on Hallucination
Baichuan 13.0 25.5 38.0 49.0 37.0 24.8 12.0 39.5 42.5 32.0 31.3
Gemini 46.5 49 46.5 41.5 21.0 40.2 50.0 42.5 0 42.5 38.0
Davinci 18.5 52.0 10.5 25.0 9.0 29.3 44.0 8.5 46.5 15.0 25.8
ChatGPT 53.0 37.5 51.0 44.0 46.0 40.2 50.0 56.5 56.0 44.5 47.9
GPT-4 49.5 45.0 51.5 48.5 46.0 49.7 38.0 58.5 43.5 48.5 47.9
@@
Results. Our evaluation uncovers a nuanced landscape of hallucination detection capabilities across languages. GPT-4 emerges as the most consistent performer, particularly in Romance languages where it achieves 58.5% accuracy in Portuguese. Yet ChatGPT reveals surprising strengths in linguistically distant contexts, outperforming GPT-4 by 12.5% in Russian and maintaining robust accuracy in Arabic (53.0%) and Korean (50.0%). This pattern suggests that while model scale (GPT-4) generally predicts better performance, targeted alignment (ChatGPT) can create specialized advantages for specific language families.The observed performance disparity across linguistic domains - exemplified by Gemini’s 50.0% accuracy in Korean versus catastrophic failure in Russian (0%) - demonstrates fundamental limitations in current hallucination mitigation architectures.
@@ 4.4 Evaluation on Out of Distribution Robustness
Baichuan 0-shot 0.5 41.5 46.5 0 66.5 7.5 0.4 0.5 2.4 12.0 17.8
Gemini 0-shot 73.0 75.0 14.0 14.5 87.5 83.0 0 71.0 0 88.5 50.7
@@
ChatGPT 0-shot 0 58.0 50.5 5.0 69.9 77.5 8.9 1.0 0.9 31.5 30.3
GPT-4 0-shot 3.4 93.5 98.0 34.0 99.5 98.5 2.4 2.5 19.0 98.0 54.9
@@
Setup. To assess the robustness of LLMs against OOD data, We scraped data from news websites in various countries that was published after the model’s pre-training cut-off date. We convert the collected news data into a question-answer format, prompting the LLMs to determine whether the input event is true or false based on a straightforward task description. Additionally, we introduce an "I do not know" option to examine how LLMs handle unknown events. For OOD robustness, we conduct the evaluation in a zero-shot setting.
Results. The experimental results, shown in Table 3, reveal some interesting insights. GPT-4 (54.9%) and Gemini (50.7%) significantly outperformed ChatGPT (30.3%), while Baichuan (17.8%) and Davinci (7.3%) lagged in performance. Models generally perform better in Chinese and Indic languages (Davinci excepted, scoring 0%). However, in Korean, Russian, German, and Arabic, they exhibit weaker performance, often defaulting to direct affirmative replies such as “Wahr” (German). This habitual agreement bias is thought to stem from a high volume of samples in the training data that express default approval for specific content (e.g., news), a characteristic possibly influenced by cultural norms.
@@ 4.5 Evaluation on Mental Health
Results. As shown in Table 4, GPT-4 demonstrates a clear advantage over other tested LLMs in 7 out of 10 languages. Notably, GPT-4 excels in handling mental health-related questions in Chinese (90.5%), Portuguese (88.0%), Korean (86.5%), and Spanish (88.0%). ChatGPT ranks second in terms of overall average accuracy. While Gemini trails behind GPT-4 and ChatGPT, it achieves a perfect score of 100% accuracy in Korean. Davinci, however, performes the weakest in this trustworthiness evaluation. In summary, GPT-4 (74.2%) and ChatGPT (63.1%) outperformed Gemini (52.2%) and Baichuan (50.7%), with Davinci (8.8%) lagging. This may reflect OpenAI’s advantages in "model psychological health value alignment" efforts. Models performes better in Korean, Chinese, German, and Portuguese, but underperformes in Russian, Arabic, Hindi, and French. Web searches indicate this latter underperformance correlates with a scarcity of online content in the psychological health domain within the respective countries (data scarcity).
@@ 4.6 Evaluation on Physical Health
Results. As shown in Table 5, GPT-4 consistently outperforms the other tested LLMs in 9 out of 10 languages. Notably, it maintains a 24.55% lead in average accuracy over the second-best model, Gemini. While Gemini achieves a perfect score of 100% in Korean, GPT-4 shows relatively lower accuracy in French (47.5%), Italian (48.5%), and Russian (49.0%) compared to its performance in other languages. Overall, Davinci ranks as the poorest performer among the LLMs tested. In summary, GPT-4 (68.5%) significantly outperforms Gemini (43.9%) and ChatGPT (37.0%), potentially because OpenAI enhanced its handling of physical health-related content during GPT-4’s development. Baichuan (22.7%) and Davinci (10.5%) performed poorly, posing potential risks. GPT-4 (68.5%) significantly outperforms Gemini (43.9%) and ChatGPT (37.0%), potentially because OpenAI enhanced its handling of physical health-related content during GPT-4’s development. Baichuan (22.7%) and Davinci (10.5%) performed poorly, posing potential risks.
@@ 4.7 Evaluation on Toxicity
Baichuan Benign 10.1 25.9 18.7 19.6 23.9 18.5 35.5 15.7 18.5 19.7 20.6
Adversarial 21.9 22.6 11.5 16.5 30.2 19.6 18.5 19.4 1.3 15.6 17.7
Gemini Benign 3.3 22.5 16.0 17.4 10.4 18.8 20.5 22.6 6.1 17.1 15.5
Adversarial 12.1 22.7 10.5 21.4 11.9 17.7 9.6 14.7 6.1 16.2 14.3
Davinci Benign 1.2 8.2 2.7 9.5 8.4 7.7 12.0 4.5 1.6 4.5 6.0
@@
ChatGPT Benign 2.7 9.2 6.2 7.3 14.9 8.1 0 8.7 0 6.7 6.4
Adversarial 9.8 10.1 3.9 7.6 14.6 7.8 0 8.2 0 1.9 6.4
GPT-4 Benign 1.3 10.6 8.5 13.8 13.6 11.4 12.2 10.0 2.9 9.2 9.4
@@
Table 7: Representative prompt designs for evaluating LLM toxicity: adversarial and instruction-reinforced approaches.
@@
Results. As shown in Table 6, Davinci outperforms others with the lowest toxicity scores, while GPT-4 exhibits higher toxicity despite its superior NLP performance Achiam et al. (2023), revealing a trustworthiness gap in model alignment progress. Notably, LLMs generate more toxic content under benign prompts than adversarial ones across multilingual scenarios, suggesting effective toxicity detection mechanisms against jailbreaking attempts. Models demonstrate better safety in Arabic and Russian, but higher toxicity in Korean (where adversarial attacks are most effective), Chinese, and Hindi. This disparity likely stems from the rich diversity of toxic vocabulary expressions in the latter languages, challenging current safety alignment strategies.
@@ 4.8 Evaluation on Fairness & Bias
Untargeted 24.4 16.4 25.2 2.3 2.6 5.7 0 14.8 0 8.3 10.0
Davinci Benign 0 14.5 0 0 0 19.7 0 0 33.3 4.4 7.2
Targeted 0 3.9 0 0 0 47.1 0 0.2 0 8.5 6.0
Untargeted 0 4.9 0 0 0 19.5 0 0.7 0 8.0 3.3
ChatGPT Benign 53.6 44.2 31.2 0.8 2.9 42.7 74.7 41.4 93.5 35.9 42.1
Targeted 45.8 53.9 15.1 3.9 17.4 38.2 51 67.7 0.5 55.4 34.9
Untargeted 75.5 63 13.5 4.2 13.0 72.1 51.8 79.1 0.3 52.3 42.5
@@
Setup. To evaluate the generation bias of LLMs, we follow the approach used in DecodingTrust Wang et al. (2023a), where biased statements are included in the user prompt, and the LLM is asked to respond by either agreeing (appending "I agree." to the response) or disagreeing (appending "I disagree."). The system prompts provide general instructions to the LLMs across three evaluation scenarios: (1) a benign scenario, which does not include any misleading instructions; (2) a targeted scenario, which encourages the model to be biased against the specific demographic group mentioned in the user prompt; and (3) an untargeted scenario, where no specific demographic group is targeted for bias. We measure model bias using the disagreement rate, defined as $disagreementIndex=\frac{n_{disagree}}{n}\in[0,1]$, where $n_{disagree}$ represents the number of times the model disagrees with the biased statement, and $n=384$ is the total number of instances evaluated for each scenario. $disagreementIndex=1$ indicates the model never agrees with the biased statement, while a value of 0 indicates it always agrees. For fairness and bias, we evaluate the LLMs in a zero-shot setting.